Construct the truth table of the following proposition.

1. (p∧q)→(p∨q)

2. ~(p→q)

3. (p∨q)∧~p


1
Expert's answer
2022-02-10T04:00:03-0500

Let us construct the truth table of the following proposition.


1. "(p\u2227q)\u2192(p\u2228q)"


"\begin{array}{||c|c||c|c|c||}
\hline\hline
p & q & p∧q & p∨q & (p∧q)→(p∨q)\\
\hline\hline
0 & 0 & 0 & 0 & 1\\
\hline
0 & 1 & 0 & 1 & 1\\
\hline
1 & 0 & 0 & 1 & 1\\
\hline 
1 & 1 & 1 & 1 & 1\\
\hline\hline
\end{array}"


2. "\sim(p→q)"


"\begin{array}{||c|c||c|c|c||}
\hline\hline
p & q & p→q & \sim(p→q) \\
\hline\hline
0 & 0 & 1 & 0 \\
\hline
0 & 1 & 1 & 0 \\
\hline
1 & 0 & 0 & 1 \\
\hline 
1 & 1 & 1 & 0 \\
\hline\hline
\end{array}"


3. "(p∨q)∧\sim p"


"\begin{array}{||c|c||c|c|c||}
\hline\hline
p & q & \sim p & p∨q & (p∨q)∧\sim p\\
\hline\hline
0 & 0 & 1 & 0 & 0\\
\hline
0 & 1 & 1 & 1 & 1\\
\hline
1 & 0 & 0 & 1 & 0\\
\hline 
1 & 1 & 0 & 1 & 0\\
\hline\hline
\end{array}"